# Can AI Prompt Humans? Multimodal Agents Prompt Players’ Game Actions and Show Consequences to Raise Sustainability Awareness

### v1

@@ 7. Limitations and Future Work
There are several limitations that should be considered in the context of our findings:
7.1. The sample’s strong sustainability awareness may bias the results.
Our sample was relatively small and predominantly comprised young individuals with higher education (undergraduate and above), a demographic that generally exhibits strong sustainability awareness (Adu et al., 2021; Lazăr et al., 2022). This may have introduced a bias in our results. Had we included participants with lower environmental awareness or those opposed to sustainability, the outcomes might have differed significantly. These participants could have shown either a notable positive shift in their environmental attitudes or, conversely, their opposition to sustainability might have resulted in no effect. In future studies, we plan to target this specific demographic to explore these potential variations.
7.2. Study duration limits assessment of lasting sustainable behavior change.
@@
7.3. Without control groups, ruling out other factors’ influence is challenging.
@@
7.4. Game elements generated by generative AI are not always satisfactory.
@@ 8. Conclusion
This study presents an attempt of leveraging the multi-modality of Generative AI to create a role-playing game and promote intended sustainability awareness. The game empowers players to interact with AI agents and convert natural conversations into concrete in-game actions with visible environmental consequences. By allowing players to experience the long-term effects of their decisions, the game fosters a deeper understanding of sustainability issues. Our findings reveal that the gameplay can motivate intended real-world behavioral changes without requiring significant shifts in players’ core beliefs or attitudes towards sustainability. This suggests that game’s design may be effective across diverse audiences, regardless of their initial sustainability awareness levels. Our research highlights the potential of GenAI-driven multimodal agents and action-consequence mechanics to create immersive gaming experiences that can inspire positive change in real-world sustainability behaviors.
9. Appendix
@@ Appendix A Background Design for Non-Player Characters (NPCs)
Character Prompts
For each NPC, a specific prompt is designed. For instance, part of Lisa’s prompt is as follows:
@@
A.1. Pre-Survey Questions and Their Types
Table 4. Pre-Survey Questions and Their Types
@@ A.1. Pre-Survey Questions and Their Types
A.2. Post-Survey Questions and Their Types
Table 5. Post-Survey Questions and Their Types

### v4

@@ 7. Limitations and Future Work
7.1. The Sample’s Sustainability Awareness May Bias The Results
Our sample was relatively small and predominantly comprised young individuals with higher education (undergraduate and above), a demographic that generally exhibits strong sustainability awareness (Adu et al., 2021; Lazăr et al., 2022). This may have introduced a bias in our results. Additionally, the demographic composition of our participants was relatively homogeneous in terms of age, education level, and geographic background, which may limit the generalizability of our findings. A more diverse participant pool, including individuals with varied environmental awareness, different cultural perspectives, and broader geographic representation, could yield different outcomes. For instance, participants with lower exposure to sustainability concepts might exhibit distinct engagement patterns or attitude shifts compared to those already inclined toward environmental consciousness.
Moreover, while we propose that the game could engage a broader (non-gaming) audience, the specific target demographic remains to be clearly defined. Future research should refine the intended audience scope and evaluate the game’s effectiveness across a more diverse range of participants to better understand its broader impact.
7.2. Impact of Social Desirability Effect and Priming Effect on Results
The evaluation of EcoEcho’s impact on intended sustainable behaviors may be affected by the social desirability effect (Grimm, 2010), where participants report socially acceptable intentions, and the behavioral priming effect (Doyen et al., 2012), where the study design influences responses aligned with its goals. These biases could overstate the game’s effectiveness. We utilized in-game assessments as implicit measures, mapping them to in-game behaviors to mitigate the influence of these two effects. Players voted on whether to support renewable energy within the game’s narrative environment, and the results were analyzed in conjunction with survey and interview data. However, in-game assessments may not fully eliminate these effects and could introduce additional influencing factors, as players’ decisions might be shaped by the game’s narrative context. In the future, employing more diverse assessment methods, such as longitudinal studies to track real-world behavior over time after gameplay, or controlled experiments comparing the gameplay group with a control group, could provide deeper insights into how gameplay influences decision-making and attitudes.
7.3. Study Duration Limits Assessment Of Lasting Sustainable Behavior Change
@@
7.4. Without Control Groups, Ruling Out Other Factors’ Influence Is Challenging
@@
7.5. Game Elements Generated By Generative AI Are Not Always Satisfactory
@@ 8. Conclusion
This study presents an attempt of leveraging the multi-modality of Generative AI to create a role-playing game and promote intended sustainability awareness. The game empowers players to interact with AI agents and convert natural conversations into concrete in-game actions with visible environmental consequences. By allowing players to experience the long-term effects of their decisions, the game fosters a deeper understanding of sustainability issues. Our findings reveal that the gameplay can motivate intended real-world behavioral changes without requiring significant shifts in players’ core beliefs or attitudes towards sustainability. Viewed more broadly, this paper demonstrates how multimodal agents and action-consequence mechanics can harness AI to effectively drive positive societal change.
@@ Appendix A Background Design for Non-Player Characters (NPCs)
Appendix B Character Prompts
@@
B.1. Pre-Survey Questions and Their Types
Table 4. Pre-Survey Questions and Their Types.
@@ B.1. Pre-Survey Questions and Their Types
B.2. Post-Survey Questions and Their Types
Table 5. Post-Survey Questions and Their Types.